OpenStudy (anonymous):

What is the tenth term of the geometric sequence 3, 6, 12, 24, 48, … ?

OpenStudy (welshfella):

nth term = a*r^(n-1) where a = first term and r = common ratio here a = 3 and r = 6/3 = 2

OpenStudy (welshfella):

10th term = 3* 2^(10-1)

OpenStudy (anonymous):

1536?

OpenStudy (welshfella):

grab your calculator and see what you get..

OpenStudy (welshfella):

correct

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw

OpenStudy (lopus):

3 = 3* (2^(1-1)) n=1 3*2 = 3* (2*(2-1)) n=2 3*2*2 = 3* (2*(3-1)) n=3 3*2*2*2 = 3* (2*(4-1)) n=4 then 3*( 2^(N-1)) if n-->1